what is the vertex of f(x)=-x^2+8x-19
Answer by stanbon(75887)

what is the vertex of f(x)=-x^2+8x-19
---
a = -1 ; b = 8 ; c = -19
---
Vertex occurs where x = -b/(2a) = -8/(2*-1) = 4
f(4) = -16 + 32 -19 = -3
---
Vertex at (4,-3)
===================
Cheers,
Stan H.
============